OpenStudy (anonymous):

Please help me.. what is the point-slope form of the equation below that represents the line that passes through the point (6, −3) and has a slope of one half..?

OpenStudy (anonymous):

start with \[y+3=\frac{1}{2}(x-6)\] and then solve for \(y\)

OpenStudy (anonymous):

you good from there?

OpenStudy (anonymous):

Yeah, i THINK i got it :)

OpenStudy (anonymous):

ok good

OpenStudy (anonymous):

you want me to check your answer?

OpenStudy (anonymous):

yes please :) just give me a second..

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

y = 1/2x − 6 ?

OpenStudy (anonymous):

good work

OpenStudy (anonymous):

thank you :)